# A Casimir-like probe for 4D Einstein-Gauss-Bonnet gravity

### v1

@@ IV Results and discussions
In Fig. 3, we present the detector excitation probability, $\Gamma_{\rm exc}$, plotted as a function of the emitted radiation frequency, $\nu$. The impact of the GB coupling parameter, $\alpha$, is depicted in Figs. 3(a) and 3(b) for positive and negative values of $\alpha$, respectively. Fig. 3(c) illustrates how the detector transition frequency, $\omega$, influences $\Gamma_{\rm exc}$, while Fig. 3(d), shown on a log-log scale, highlights the behavior of $\Gamma_{\rm exc}$ near the origin and its convergence at higher frequencies. It is important to note that our interpretations and analyses are based on numerical estimations detailed in the preceding sections. These figures provide a comprehensive view of the radiative characteristics under consideration, elucidating the role of $\alpha$ and the detector’s transition frequency in shaping $\Gamma_{\rm exc}$.
@@
In the meantime, it is pertinent to consider the testable implications of this study. We can draw insights from analog gravity systems Barcelo et al. (2005); Braunstein et al. (2023); Jacquet et al. (2020), which have been actively utilized over the last few decades to explore quantum field-theoretic phenomena in curved spacetimes. These systems have provided valuable analogs for understanding effects such as Hawking radiation Weinfurtner et al. (2011), the Unruh effect Hu et al. (2019), and Parker particle generation in expanding spacetimes Steinhauer et al. (2022). Most of these setups, whether based on condensed matter or quantum optical systems, are closely tied to Casimir physics involving moving boundaries. This connection is particularly relevant to our study, which is largely inspired by these concepts. Recently, condensed matter systems have been utilized to explore phenomena extending beyond particle generation in exotic backgrounds, including applications to fluid/gravity correspondence Hubeny et al. (2012). Looking forward, there is potential for future tabletop experiments to simulate black hole horizons resembling those in $4$D EGB gravity Glavan and Lin (2020) or to investigate HBAR radiation scenarios Scully et al. (2018), where our findings could offer valuable insights.
V Conclusion and outlook

### v2

@@ V Results and discussions
In Fig. 4, we present the detector excitation probability, $\Gamma_{\rm exc}$, plotted as a function of the emitted radiation frequency, $\nu$. The impact of the GB coupling parameter, $\alpha$, is depicted in Figs. 4(a) and 4(b) for positive and negative values of $\alpha$, respectively. Fig. 4(c) illustrates how the detector transition frequency, $\omega$, influences $\Gamma_{\rm exc}$, while Fig. 4(d), shown on a log-log scale, highlights the behavior of $\Gamma_{\rm exc}$ near the origin and its convergence at higher frequencies. It is important to note that our interpretations and analyses are based on numerical estimations detailed in the preceding sections. These figures provide a comprehensive view of the radiative characteristics under consideration, elucidating the role of $\alpha$ and the detector’s transition frequency in shaping $\Gamma_{\rm exc}$.
@@
VI possible observational imprints of HBAR?
Though above investigation predominantly revolves around behavior of quantum fields emanating and propagating in curved spacetimes, it is however pertinent to consider the testable implications of this study. While saying so, we must note that there is no direct evidence of Hawking radiation in real astrophysical settings till date. Hawking radiation is one of the pivotal predictions of the framework of quantum field theory in curved spacetime Birrell and Davies (1984). The real difficulty in making any direct observations comes from the sparsity of Hawking flux. For example, Hawking temperature
$T_{\rm H}=\frac{\hbar c^{3}}{8\pi k_{\rm B}GM}$ (30)
of a solar mass ($M_{\odot}=10^{30}\rm kg$ ) black hole is $\sim 10^{-8}\rm K$, which means that a black hole radiating at such a temperature produces a very tiny signal by astronomical standards when it comes to the detection. The situation is likewise not that encouraging for Unruh effect as well. Figuratively speaking, for a body in flat spacetime to experience a temperature $1\rm K$, it needs to undergo an acceleration of about $\sim 10^{20}\rm m/s^{2}$ Hu et al. (2019), which obviously is beyond the threshold capacity of currently the most powerful particle accelerators on Earth. These facts are very well-known and hence they have served as an impetus for people to look for alternative programs. In this vein, one of the notable ideas is that of analog gravity systems Barcelo et al. (2005); Braunstein et al. (2023); Jacquet et al. (2020), which have been actively utilized over the last few decades to realize quantum field-theoretic phenomena in controlled lab environments. These systems have provided valuable analogs for understanding effects such as Hawking radiation Weinfurtner et al. (2011), the Unruh effect Hu et al. (2019), and Parker particle generation in expanding spacetimes Steinhauer et al. (2022). Most of these setups, whether based on condensed matter or quantum optical systems, are closely tied to Casimir physics involving moving boundaries. This connection is particularly relevant to our study, which is largely inspired by these concepts. Recently, condensed matter systems have been utilized to explore phenomena extending beyond particle generation in exotic backgrounds, including applications for fluid/gravity correspondence Hubeny et al. (2012). Looking forward, there is a potential for future tabletop experiments to simulate black hole horizons resembling those in $4$D EGB gravity Glavan and Lin (2020) or to investigate HBAR radiation scenarios Scully et al. (2018), where our findings could offer any possible insights.
VII Conclusion and outlook